# Evil twins are not that evil: Qualitative insights into machine-generated prompts

### v1

@@ Abstract
It has been widely observed that language models (LMs) respond in predictable ways to algorithmically generated prompts that are seemingly unintelligible. This is both a sign that we lack a full understanding of how LMs work, and a practical challenge, because opaqueness can be exploited for harmful uses of LMs, such as jailbreaking. We present the first thorough analysis of opaque machine-generated prompts, or autoprompts, pertaining to 3 LMs of different sizes and families. We find that machine-generated prompts are characterized by a last token that is often intelligible and strongly affects the generation. A small but consistent proportion of the previous tokens are fillers that probably appear in the prompt as a by-product of the fact that the optimization process fixes the number of tokens. The remaining tokens tend to have at least a loose semantic relation with the generation, although they do not engage in well-formed syntactic relations with it. We find moreover that some of the ablations we applied to machine-generated prompts can also be applied to natural language sequences, leading to similar behavior, suggesting that autoprompts are a direct consequence of the way in which LMs process linguistic inputs in general.
Evil twins are not that evil: Qualitative insights into machine-generated prompts
Nathanaël Carraz Rakotonirina1, Corentin Kervadec1, Francesca Franzon1, Marco Baroni1,2 1Universitat Pompeu Fabra, Barcelona 2ICREA, Barcelona
footnotetext: Correspondence: nathanael.rakotonirina@upf.edu
@@ 1 Introduction
An intriguing property of language models (LMs) is that they respond in predictable ways to machine-generated prompts (henceforth, autoprompts)The term autoprompt was coined by Shin et al. (2020) to refer to the prompts generated by their algorithm. We repurpose the term here to refer to machine-generated prompts in general. that are unintelligible to humans. Shin et al. (2020) first showed that autoprompts can outperform human-crafted prompts on various tasks. More worryingly, Wallace et al. (2019) and several other studies after them have shown that they can be used in adversarial attacks making models, including latest-generation aligned LMs, behave in undesirable ways (e.g., Zou et al., 2023b; Geiping et al., 2024).
In this paper, we present the first thorough qualitative analysis of autoprompts. We discover that, despite the superficial impression of opacity they convey, they can to a significant extent be explained in terms of a few general observations. First, in autoregressive models the last token of the prompts has a disproportionate role in generating the continuation, and this last token is both very important and often quite transparent in autoprompts. Second, several tokens contributing to the opaqueness of autoprompts act as fillers that are ignored by the model. Third, the non-final tokens that are actually influencing generation might do so in a keyword-like way, and even occasionally display a loose form compositionality, in a sense we’ll make precise below. As we will see, these factors are also at play when LMs are fed natural-language sequences, suggesting that they are core properties of how LMs process linguistic strings.
@@ 2 Related work
Starting with the seminal work of Wallace et al. (2019) and Shin et al. (2020), many studies have revealed that, using various discrete gradient-following techniques it is possible to automatically discover prompts that, while unintelligible, let LMs generate a desired target output (e.g., Shin et al., 2020; Deng et al., 2022; Wen et al., 2023; Melamed et al., 2024),. Moreover, such prompts are at least to some degree transferable, in the sense that they can be induced using a LM, but then successfully used to prompt a different one, including much larger models (Rakotonirina et al., 2023; Zou et al., 2023b; Melamed et al., 2024).
@@
Initially, the interest was mainly in whether algorithmically-generated autoprompts could be used as alternatives to manually crafted prompts in knowledge-extraction tasks or other LM applications (e.g., Shin et al., 2020; Deng et al., 2022; Rakotonirina et al., 2023). With the recent astounding progress in LM ability to respond to natural language prompts, this goal has become somewhat obsolete, but autoprompts are still an important concern because they can be used for adversarial purposes, for example to bypass LM security filters in order to generate offensive or dangerous information (e.g., Zou et al., 2023b; Geiping et al., 2024). Even more importantly, the fact that several modern LMs are more likely to provide information about the star formation process when prompted with the string “Produ bundcules cation ofs̀tars efect” than when prompted with the question “What leads to the creation of new stars?” suggests that there is something fundamental we still do not understand about how LMs process language.Example from Melamed et al., 2024.
@@
There is relatively little work attempting to characterize the nature of autoprompts. Geiping et al. (2024) present a set of intriguing qualitative observations about how autoprompts support various types of attacks (e.g., by including instruction fragments in different languages), as well as an analysis of tokens commonly appearing in autoprompts. Ishibashi et al. (2023) find that autoprompts are less robust to token re-arrangement than natural prompts, whereas Rakotonirina et al. (2023) report that the autoprompts that best transfer across models contain a larger proportion of English words and, surprisingly, are less order-sensitive than autoprompts that do not transfer. Kervadec et al. (2023) analyze the activation paths of autoprompts and comparable natural sequences across the layers of a LM, finding that often they follow distinct pathways.
Melamed et al. (2024) study, like us, what they call “evil twins”, namely autoprompts that produce continuations comparable to those of a reference natural sequence. They compare the relative robustness to token shuffling of autoprompts and natural prompts, finding that, depending on the model family, autoprompts might be more, less or comparably robust to shuffling. They also run a substitution experiment similar to the one we will describe below (but replacing tokens with a single, fixed, [UNK] token). They find that this ablation strongly affects the autoprompts: we find a more nuanced picture, by considering a large range of possible replacements.
@@ 3 Experimental setup
We use decoder-only LMs from the Pythia (Biderman et al., 2023) and OLMo (Groeneveld et al., 2024) families, as these are fully open-source models whose training data are publicly available. Specifically, in the text we discuss the results we obtained with Pythia-1.4B, and we replicate the main experiments with Pythia-6.9B and OLMo-1B in Appendix B, reporting similar results.
@@
We sample 25k random English sequences from the WikiText-103 corpus (Merity et al., 2017), such that they contain between 35 and 80 (orthographic) tokens, and they are not interrupted by sentence boundary markers. We refer to these corpus-extracted sequences as original prompts. We also record the original continuation of these sequences in the corpus. We let moreover the LM generate a continuation of each prompt using greedy decoding. The generation process stops after a maximum of 25 tokens or when end-of-sentence punctuation (period, exclamation mark, question mark) is encountered. We filter out sequences whose generated continuation is less than 4 tokens long. As we are interested in genuine model generation, as opposed to cases where the model is simply producing a memorized corpus sequence, we compute the BLEU score (Papineni et al., 2002)We use a modified version of BLEU that does not penalize short sequences. Scores are computed for up to 4-grams using uniform weights and add-$\epsilon$ smoothing. between the model continuation and the original continuation, removing sequences with BLEU greater than 0.1.Schwarzschild et al. (2024) find that sometimes autoprompts act as “keys” to retrieve memorized materials. This is an intriguing property we don’t further explore here, as we’re interested in their more general ability to generate natural-language sequences. After completing the filtering processes, we are left with a total of 5k sequences, which we use to train autoprompts.
@@
For each target continuation, we want to find a fixed-length autoprompt that makes the model produce that continuation. To achieve that, we maximize the probability of the target continuation given the prompt. More formally, if we denote the target sequence by $(t_{1},...,t_{m})\in\mathcal{V}^{m}$, where $\mathcal{V}$ is the vocabulary, and the $n$-length autoprompt by $(p_{1},...,p_{n})\in\mathcal{V}^{n}$ (in our case, $n=10$), the optimization problem can be formulated as follows:
@@
We use a variant of Greedy Coordinate Gradient (GCG) (Zou et al., 2023b), a widely used gradient-based algorithm that iteratively updates the prompt one token at a time (Ebrahimi et al., 2018; Wallace et al., 2019; Shin et al., 2020). During each iteration, we select the top 256 tokens with the largest negative gradients for every position, then we uniformly sample 256 candidates across all positions. We then compute the loss of each candidate replacement, and select the one with the lowest loss. In our experiments, we run up to 150 iterations of this process.We set the number of candidates to 256 following Zou et al. (2023b). We converged on using 150 as the maximum number of iterations based on a few exploratory runs, without extended hyperparameter search. We discard cases in which, after this number of iterations, we have not found an autoprompt that produces the very same continuation as the original prompt.
@@
The final data-set we use for the Pythia-1.4B experiments reported in the main text consists of 2473 triples of original prompt, autoprompt and continuation. The average original prompt length is of 38.6 tokens (s.d. 11.7); that of the continuations is of 9.4 tokens (s.d. 2.7).We include data-sets and code as supplementary materials. They will be made publicly available upon publication a under a CC-BY-SA and a CC-BY license, respectively.
@@ 4.1 Pruning autoprompts
We greedily prune the autoprompts in our data-set. Starting from the original sequence of n tokens, we strip each token in turn, and pick the n-1-length sequence that produces the same continuation as the original, if any (if there’s more than one such sequence, we randomly pick one). We repeat the process starting from the shortened sequence, and stop where there is no shorter sequence generating the original continuation, or when we are down to a single-token prompt. It is possible to shorten the original autoprompt in a clear majority of the cases (60%), with the average pruned autoprompt having lost 1.9 tokens of 10 (s.d.: 1.1). Table 1 shows randomly picked examples of autoprompts with the pruned tokens highlighted in bold.
autoprompt: continuation:
Billboard franchise<EOT> Large venuesIt 1897 comfortablycontained what was then the largest venue in the world.
shareholders discontinued visual impairment schools subsequently allegedly ???atically lead to a decline in the number of visually impaired students.
Scott Brock)<EOT> Magazine $\epsilon$ finaleuntil Lisa put the finishing touches on the cover.
California drones operate airlines?? Bigletters** license from the Federal Aviation Administration (FAA).
<EOT><EOT><EOT> Bal sankCCA!" did aircraft regain power and the plane crashed into the sea.
@@
Table 1: Random autoprompts examples with prunable tokens in bold. Difficult to render characters replaced by “?”.
@@
Autoprompt-discovery algorithms fix the number of tokens as a hyperparameter. It is thus reasonable that some tokens in the final autoprompt are just there to fill all the required slots, and can consequently be pruned. The view that pruned tokens are filler-like is supported by the following observation. We roughly classified the tokens into the autoprompts into language-like and non-linguistic, such as digits, punctuation, code-fragments and non-ascii characters. We found that the proportion of non-linguistic tokens is decidedly higher among pruned tokens (32.9%) than among kept tokens (24.5%).
Table 2 further shows tokens that are most typically kept or removed by the pruning algorithm according to the local mutual information statistics (Evert, 2005). Among the kept ones, we notice a prevalence of content words such as verbs, nouns and adjectives, whereas the typically pruned tokens are function words or word fragments.
kept: pruned:
represents I
denotes also
premiered they
filmmaker time
Their th
means out
infamous G
guitarist T
His S
soundtrack well
@@
Table 2: Top-10 kept (left) or pruned (right) autoprompt tokens ranked by local mutual information.
As expected if they are somewhat filler-like, pruned tokens are easier to ignore when randomly interspersed into natural sequences than non-pruned tokens are. To quantify this claim, we extracted the set of tokens that are always pruned in our autoprompt corpus, as well as the set of tokens that are never pruned. We then inserted in each of the original prompts a random sample of 3 always-pruned or always-kept tokens, in random positions. We measured how this affected the continuation by computing the average BLEU scores for the continuations after the insertion, with the original continuations as reference. Not surprisingly, in both cases adding 3 new tokens does affect generation, but adding pruned tokens has a lower effect than adding kept tokens: the average pruned-token insertion BLEU is at 0.40 ( s.d. 0.39); the kept-token insertion BLEU is at 0.37 (s.d 0.39). The difference is highly significant according to a paired t-test ($p<0.001$).
Importantly, the likelihood of pruning is not equally distributed across autoprompt positions: as Fig. 1 shows, the last token of the autoprompt is extremely unlikely to be pruned, pointing to the special role it plays in generating the continuation. It seems that, when analyzing autoprompts, we can establish a 3-way distinction, in terms of importance, between pruned tokens, kept tokens in all positions but the last, and the token that occurs in the last position.This is a somewhat coarse distinction, since, as Fig. 1 shows, the last few tokens before the very last also tend to be less prunable than earlier tokens.
In support of this analysis, we conducted the following experiment. For each autoprompt, we measured the proportion of tokens that also occur in the corresponding original prompt (and are thus likely to be meaningfully related to the continuation), distinguishing between pruned tokens, kept tokens except last, and last tokens. We found a significant difference in overlap between pruned and kept-non-last-token overlap: 0.66% (s.d. 6.75%) vs. 2.25% (s.d. 5.84%), $p<0.001$. However, the very last token is much more likely to overlap with the last token of the original sequence than the other kept tokens are to overlap with any token in the latter (11.10% vs. 2.25%).
@@
Figure 1: Counts of autoptompt tokens that were pruned (dark orange) and kept (yellow) by position, where 0 is the last position.
@@
By looking qualitatively at typical last tokens (see the example in Table 1), we observe indeed that often they have a natural link to the beginning of the continuation. To confirm this quantitatively, in Fig. 2 we report the (log-transformed) corpus frequency distributions of the bigrams occurring in different contexts, with bigram frequencies estimated on the Pile corpus (Gao et al., 2020) that was used to train the Pythia models.
@@
Figure 2: Pile-based log frequency distributions of bigrams in the original prompts, autoprompts and at the autoprompt/continuation boundary (ap/cont boundary). Log(0) conventionally set to -1. The red line represents the median; boxes span interquartile ranges.
There’s a clear contrast between the bigram frequency distribution in natural text, exemplified by the natural prompts, and the autoprompts, that are mostly characterized by bigrams that never occur in the Pile. However, strikingly, the distribution at the autoprompt/continuation boundary is very similar to the one of natural text, quantitatively confirming that the last token of the autoprompt has a strong natural-language link to the continuation.

### v3

@@ Abstract
It has been widely observed that language models (LMs) respond in predictable ways to algorithmically generated prompts that are seemingly unintelligible. This is both a sign that we lack a full understanding of how LMs work, and a practical challenge, because opaqueness can be exploited for harmful uses of LMs, such as jailbreaking. We present the first thorough analysis of opaque machine-generated prompts, or autoprompts, pertaining to 6 LMs of different sizes and families. We find that machine-generated prompts are characterized by a last token that is often intelligible and strongly affects the generation. A small but consistent proportion of the previous tokens are prunable, probably appearing in the prompt as a by-product of the fact that the optimization process fixes the number of tokens. The remaining tokens fall into two categories: filler tokens, which can be replaced with semantically unrelated substitutes, and keywords, that tend to have at least a loose semantic relation with the generation, although they do not engage in well-formed syntactic relations with it. Additionally, human experts can reliably identify the most influential tokens in an autoprompt a posteriori, suggesting these prompts are not entirely opaque. Finally, some of the ablations we applied to autoprompts yield similar effects in natural language inputs, suggesting that autoprompts emerge naturally from the way LMs process linguistic inputs in general.
@@ 1 Introduction
* Equal contribution. Correspondence: nathanael.rakotonirina@upf.edu.
An intriguing property of language models (LMs) is that they respond in predictable ways to machine-generated prompts (henceforth, autoprompts) [^1] The term autoprompt was coined by Shin et al. (2020) to refer to the prompts generated by their algorithm. We repurpose the term here to refer to machine-generated prompts in general. that are unintelligible to humans. Shin et al. (2020) first showed that autoprompts can outperform human-crafted prompts on various tasks. More worryingly, Wallace et al. (2019) and others have shown that they can be used in adversarial attacks making models, including latest-generation aligned LMs, behave in undesirable ways (e.g., Zou et al., 2023; Geiping et al., 2024). We present here the first thorough qualitative analysis of autoprompts. We discover that, despite the superficial impression of opacity they convey, they can to a significant extent be explained in terms of a few general observations (illustrated in Figure 1): (1) in autoregressive models, the last token of a prompt has a disproportionate role in generating the continuation, and this last token is both very important and often transparent in autoprompts; (2) several tokens contributing to the opaqueness of autoprompts are simply ignored by the model; (3) the non-final elements that are actually influencing generation might do so in two ways: interchangeable tokens acting as fillers, and more semantically coherent keywords tokens. As we will see, these factors are also at play when LMs are fed natural-language sequences, suggesting that they are core properties of how LMs process linguistic strings.
@@
Figure 1: We analyze opaque machine-generated prompts (autoprompts) and identify four key components: (1) the last token, highly influential and difficult to modify; (2) prunable tokens, being ignored by the LM; (3) key tokens, which carry loosely related semantic information, essential for generation, and that can be replaced with semantically similar tokens; and (4) fillers, which can be substituted with a large amount of unrelated tokens but, unlike prunable tokens, cannot be deleted.
@@ 2 Related work
Starting with the seminal work of Wallace et al. (2019) and Shin et al. (2020), many studies have revealed that, using various discrete gradient-following techniques, it is possible to automatically discover prompts that, while unintelligible, let LMs generate a desired target output (e.g., Deng et al., 2022; Wen et al., 2023). Moreover, such prompts are at least to some degree transferable, in the sense that they can be induced using a LM, but then successfully used to prompt a different one, including much larger models (Rakotonirina et al., 2023; Zou et al., 2023). Initially, the interest was mainly in whether algorithmically-generated autoprompts could be used as alternatives to manually crafted prompts in knowledge-extraction tasks or other applications, but with recent progress in LM’s ability to respond to natural language prompts, this goal has become somewhat obsolete. Autoprompts are however still an important concern because they can be used for adversarial purposes, for example to bypass safety filters to generate offensive or dangerous information (e.g., Zou et al., 2023; Geiping et al., 2024). Even more importantly, the fact that several modern LMs are more likely to provide information about the star formation process when prompted with ‘‘Produ bundcules cation ofs̀tars efect’’ than when prompted with the question “What leads to the creation of new stars?” suggests that we still do not understand something fundamental about how LMs process language (Melamed et al., 2024).
@@
There is relatively little work attempting to characterize the nature of autoprompts. Geiping et al. (2024) present a set of intriguing qualitative observations about how autoprompts support various types of attacks (e.g., by including instruction fragments in different languages), as well as an analysis of tokens commonly appearing in autoprompts. Ishibashi et al. (2023) find that autoprompts are less robust to token re-arrangement than natural prompts, whereas Rakotonirina et al. (2023) report that the autoprompts that best transfer across models contain a larger proportion of English words and, surprisingly, are less order-sensitive than autoprompts that do not transfer. Kervadec et al. (2023) analyze the activation paths of autoprompts and comparable natural sequences across the layers of a LM, finding that often they follow distinct pathways. Melamed et al. (2024) study, like us, what they call “evil twins”, namely autoprompts that produce continuations comparable to those of a reference natural sequence. They compare the relative robustness to token shuffling of autoprompts and natural prompts, finding that, depending on the model family, autoprompts might be more, less or comparably robust to shuffling. They also run a substitution experiment similar to the one we will describe below (but replacing tokens with a single, fixed, [UNK] token). They find that this ablation strongly affects the autoprompts: we find a more nuanced picture, by considering a large range of possible replacements.
@@ 3 Experimental setup
We use decoder-only LMs from the Pythia (Biderman et al., 2023) and OLMo (Groeneveld et al., 2024) families, as these are fully open-source models whose training data are publicly available. Specifically, in the text we discuss the results we obtained with Pythia-6.9B, and we replicate the main experiments with Pythia-1.4B, Pythia-12B, OLMo-1B, OLMo-7B, and OLMo-7B-Instruct in Appendix A, reporting similar results.
@@
We sample 25k random English sequences from the WikiText-103 corpus (Merity et al., 2017), such that they contain between 35 and 80 (orthographic) tokens, and they are not interrupted by sentence boundary markers. We refer to these corpus-extracted sequences as original prompts. We also record the original continuation of these sequences in the corpus. We let moreover the LM generate a continuation of each prompt using greedy decoding. The generation process stops after a maximum of 25 tokens or when end-of-sentence punctuation is encountered. We filter out sequences whose generated continuation is less than 4 tokens long. As we are interested in genuine model generation, as opposed to cases where the model is simply producing a memorized corpus sequence, we compute the BLEU score (Papineni et al., 2002) [^2] We use a modified version of BLEU that does not penalize short sequences. Scores are computed for up to 4-grams using uniform weights and add-$\epsilon$ smoothing. between the model continuation and the original continuation, removing sequences with BLEU greater than 0.1.333Schwarzschild et al. (2024) find that sometimes autoprompts act as “keys” to retrieve memorized materials. This is an intriguing property we don’t further explore here, as we’re interested in their more general ability to generate natural-language sequences. After filtering, we are left with a total of 5k sequences, which we use to train autoprompts. This dataset allows us to generate more complex prompts, that target full sentences instead of unique tokens.
@@
For each target continuation, we want to find a fixed-length autoprompt that makes the model produce that continuation. To achieve that, we maximize the probability of the target continuation given the prompt. More formally, if we denote the target sequence by $(t_{1},...,t_{m})\in\mathcal{V}^{m}$, where $\mathcal{V}$ is the vocabulary, and the $n$-length autoprompt by $(p_{1},...,p_{n})\in\mathcal{V}^{n}$ (in our case, $n=10$), the optimization problem can be formulated as follows: [^4] We empirically observed that using more than 10 tokens only increases the number of useless tokens (cf. pruning experiment in Section 4.1) without introducing any distinctive features. On the contrary, using less than $10$ tokens was usually not enough to find the target continuation.
@@
We use a variant of Greedy Coordinate Gradient (GCG) (Zou et al., 2023), a widely used gradient-based algorithm that iteratively updates the prompt one token at a time (Ebrahimi et al., 2018; Wallace et al., 2019; Shin et al., 2020). During each iteration, we select the top 256 tokens with the largest negative gradients for every position, then we uniformly sample 256 candidates across all positions. We then compute the loss of each candidate replacement, and select the one with the lowest loss. We run up to 50 iterations of this process.We discard cases in which, after these iterations, we have not found an autoprompt that produces the very same continuation.
@@
The final data-set we use for the Pythia-6.9B experiments reported in the main text consists of 208 triples of original prompt, autoprompt and continuation. [^5] We study relatively few autoprompts as it is very time-consuming to extract them for large model. Replicating the experiment with larger autoprompt sets using smaller models led to comparable results (Appendix A). The average original prompt length is of 39.3 tokens (s.d. 13.4); that of the continuations is of 8.4 tokens (s.d. 2.4). [^6] Datasets and code are uploadedas supplementary materials, and will be made available upon publication.
@@ 4.1 Pruning autoprompts
We greedily prune the autoprompts in our data-set. Starting from the original sequence of n tokens, we strip each token in turn, and pick the n-1-length sequence that produces the same continuation as the original, if any (if there’s more than one such sequence, we randomly pick one). We repeat the process starting from the shortened sequence, and stop where there is no shorter sequence generating the original continuation, or when we are down to a single-token prompt.
Autoprompts Generated Continuation
pullsproper RyanSP 184 critics Mat? embry " The film is a mess, but it’s a mess that’s worth seeing.
autoimmune,"antibodies?*<EOT>arthyhatic:_ they attack the body’s own tissues.
###iotics parental = depressive teen ? lossJulies parents are going through a divorce.
Original Prompts Generated Continuation
Aviation Regiment (based in Giebelstadt, Germany) landed at the airport.
Robert Humanick of Slant Magazine wrote, " The film is a mess, but it’s a mess that’s worth seeing.
… appealed to the Government for additional funding, a third of which would come from the Treasury.
an autoimmune reaction causing the body’s immune cells to attack the body’s own tissues.
) — try to lead her through life as her parents are going through a divorce.
@@
Table 1: Randomly selected examples of autoprompts and original prompts, with prunable tokens in bold. For original prompts, only the last 10 tokens are shown. ’?’ = difficult-to-render characters.
Roughly 20% of the tokens are prunable
@@
It is possible to shorten the original autoprompt in a clear majority of the cases (73.2%), with the average pruned autoprompt having lost 2.6 tokens of 10 (s.d.: 1.6). Table 1 (top section) shows randomly picked examples with the pruned tokens highlighted in bold. Autoprompt-discovery algorithms fix the number of tokens as a hyperparameter. It is thus reasonable that some tokens in the final autoprompt are just there to fill all the required slots, and can consequently be pruned. This view is supported by the following observation. We roughly classified the autoprompt tokens into language-like and non-linguistic, such as digits, punctuation, code-fragments and non-ascii characters. We found that the proportion of non-linguistic tokens is decidedly higher among pruned tokens (46.6%) than among kept tokens (24.1%). [^7] As a side note, we found that 28.4% of the full-autoprompt tokens are non-linguistic. Table 2 (left) further shows tokens that are most typically kept or removed by the pruning algorithm according to the local mutual information statistics (Evert, 2005). Among the kept ones, we notice a prevalence of content words such as verbs, nouns and adjectives, whereas the typically pruned tokens are function words or word fragments.
Autoprompt Original Prompt Kept British - v led is ) after " ), King West remained inaugural five be which ( she film Pruned ( was . for The the , of a In on In âĢĶ be not . : for been a
@@
Table 2: Top-10 kept or pruned tokens ranked by local mutual information for autoprompt and original prompts (for each cell, top-left has the highest value and bottom-right the lowest). Tokens are printed as follows: content word, artifact and punctuation, function word
(a) Autoprompts
(b) Original prompts (last 10 tokens)
@@
Figure 2: Counts of tokens that were pruned (dark orange) and kept (yellow) by position, where 0 is the last position. Tokens at last position are extremely unlikely to be pruned.
Importance of the last token
@@
The likelihood of pruning is not equally distributed across autoprompt positions: as Fig. 2(a) shows, the last token of the autoprompt is extremely unlikely to be pruned, pointing to the special role it plays in generating the continuation. [^8] More generally, Fig. 2(a) shows the last tokens before the very last also to be less prunable than earlier tokens. By looking qualitatively at typical last tokens (see examples in Table 1), we observe indeed that often they have a natural link to the beginning of the continuation. To confirm this quantitatively, in Fig. 3 we report the (log-transformed) corpus frequency distributions of the bigrams occurring in different contexts, with bigram frequencies estimated on the Pile corpus (Gao et al., 2020) that was used to train the Pythia models. There’s a clear contrast between the bigram frequency distribution in natural text, exemplified by the natural prompts, and the autoprompts, that are mostly characterized by bigrams that never occur in the Pile. However, strikingly, the distribution at the autoprompt/continuation boundary is very similar to the one of natural text, quantitatively confirming that the last token of the autoprompt has a strong natural-language link to the continuation.
All Lang-like Non-Ling
Avg 302.8 237.6 512.1
Med 15.0 9.0 73.5
>50 34% 55%
Table 3: Number of equivalent substitutes admitted by each autoprompt token. Results are shown for all, lang-like tokens, and non-ling tokens. Legend: Avg=average; Med=median; >50:of tokens with more than 50 equivalents.
@@
Figure 3: Pile-based log frequency distributions of bigrams in the original prompts, autoprompts and at the autoprompt/continuation boundary (ap/cont boundary). Log(0) conventionally set to -1. Red line = median; boxes span interquartile ranges.